# Dark Dimension and the Effective Field Theory limit

### v3

@@ Abstract
Abstract
In [1] we pointed out that in the Dark Dimension scenario [2] theoretical issues arise when the prediction for the vacuum energy $\rho$, that is obtained from swampland conjectures in string theory, is confronted with the corresponding result for $\rho$ in the field theory limit. One of the problems arises from the field theory side and is due to previous results for $\rho$ that, according to our analysis, turn out to be incorrect, although widely spread. Our work was challenged in [3]. Here we show why in our opinion the claims in [3] are inconsistent, actually based on arguments deprived of physical meaning, providing further support to our findings. We conclude by discussing ideas on the physical mechanism that could trigger the suppression of large UV contributions to $\rho$.
In search for a solution to the naturalness/hierarchy problem, in the nineties the possibility of having large extra dimensions (millimetre size) was considered [5, 4]. Recently, these ideas were applied to the so-called dark dimension (DD) scenario [2], according to which we might well live in a $5D$ universe with one compact dimension of micrometre size. In a recent paper [1], we analysed aspects of this scenario, according to which above the neutrino scale our universe is described by a $(4+1)$D theory with one compact dimension of micrometer size, and pointed out possible theoretical issues. Our work was challenged in [3]. Here we show why in our opinion the authors of [3] (in their attempt to confute our results) bring confusion on the whole subject, and use incorrect arguments that lead them to physically meaningless conclusions. In view of the great wealth of theoretical and phenomenological work triggered by the DD scenario, we believe it worth trying to bring clarity on the subject and continuing its investigation.
To this end, we begin by recalling the main points [^1] For convenience of the reader less familiar with the DD proposal, in Appendix A we provide a more detailed summary of the different aspects and steps that lead to this scenario. of the DD scenario [2] together with the main results of our work [1]. In a nutshell, the DD proposal combines swampland conjectures with observational data and phenomenological bounds, and is based on the assumption that our universe lies in a unique (asymptotic) corner of the quantum gravity landscape, where only two kinds of light tower states are expected: towers of string excitations or towers of KK states. Indicating with $\mu_{tow}$ the tower scale, and with $\Lambda_{cc}$ the measured vacuum energy (cosmological constant times $M_{P}^{2}$, with $M_{P}$ the Planck mass), the application of the distance conjecture to (A)dS vacua relates the two as
@@
The authors of [3] start their rebuttal of our work [1] claiming that we question the basic (swampland) relation (4) of the DD proposal. Before going into more details, we stress that in [1] we do not question (4), nor was this our goal. What we did is different. We now briefly recall and clarify the scope and the main results of [1].
As already said, according to the proponents of the DD scenario, above the $m_{{}_{\rm KK}}$ scale our universe is described by a $(4+1)$D EFT with one compact dimension of micrometer size. This proposal has triggered a large amount of work, where $(4+1)$D EFT models that could realize the DD scenario are proposed, and their phenomenological consequences studied (see for instance [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]). In [1], our goal was to investigate on the matching between the string theory framework in which the basic relation (4) of the DD proposal is obtained and its low energy EFT limit.
@@
with no need for fine-tuning. In [1] we showed that the right hand side of (5) also contains UV-sensitive terms (see the first line in the right hand side of Eq. (Dark Dimension and the Effective Field Theory limit) below), that were missed in previous literature [^3] The presence of similar UV-sensitive terms in the calculation of the effective potential of a scalar theory was first hinted in [27]. Moreover, it was shown in [28] that the presence of Fayet-Iliopulos terms induces further divergences.. Therefore, the matching between the prediction (4) for the vacuum energy $\Lambda_{cc}$, that comes from swampland conjectures in string theory, and the corresponding result in the low energy field theory limit is not as straightforward as it was thought on the basis of the result (5). In this respect, it is worth to stress again (see the paragraph below Eq. (3)) that such a matching is crucial for the DD scenario.
In [1] we show why it is commonly thought that $\rho^{1l}$ is finite and proportional to $m_{{}_{\rm KK}}^{4}$: this result comes from a mistreatment of the loop momentum asymptotics. Performing the calculation properly, we found that (in addition to the finite term $\sim m^{4}_{{}_{\rm KK}}$ and to divergences that cancel out in SUSY theories) the one-loop contribution to the vacuum energy $\rho^{1l}$ contains “UV-sensitive” terms proportional to powers of the boundary charges. These divergences do not disappear, not even in a SUSY theory. This is due to the fact that, to trigger the Scherk-Schwarz mechanism, the boundary charges $q_{b}$ and $q_{f}$ of the boson and fermion superpartners need to be different. Below we show how the authors of [3] misuse and distort our findings, and how this leads them to meaningless conclusions.
@@
This result is important for our following analysis. Let us also note that $m_{{}_{\rm KK}}$ is the cutoff of the “reduced 4D theory”, which besides brane fields contains the zero modes of bulk fields. In Fig. 1 we provide a comprehensive picture of the general framework in which the DD scenario is framed. Finally, it is worth to observe that, although both the Standard Model and this “reduced theory” are $4$-dimensional EFTs, they are certainly not the same theory, and their cutoffs ($\Lambda_{\rm SM}$ and $m_{{}_{\rm KK}}$ respectively) should not be confused.
In Eq. (Dark Dimension and the Effective Field Theory limit), the divergences (i.e. UV-sensitive terms) are in the first line, the finite term in the second one. Limiting ourselves to consider only the most divergent term of $\rho^{1l}$ (and doing the same for the non-SUSY theory, whose result for $\rho^{1l}$ is not reported above but can be found in our work [1]), we have
@@
Although this is opposite to what we did (we repeat ourselves, in [1] our goal was to point out the presence of unexpected UV-sensitive terms), the replacement (11) is not in itself a problem. It is a theoretically legitimate possibility that the cutoff $\Lambda$ is a physical scale that appears in physical quantities, with no need for fine-tuning. However, the consequences of such a choice have to be fully taken into account. Anticipating on the results, we will show that the replacement $\rho^{1l}\to\Lambda_{cc}$ has crucial physical consequences that the authors of [3] failed to recognize. It is precisely this replacement that leads them to meaningless conclusions. Let us note that, for reasons that will be soon apparent, the right hand side of (10)1 should contain all the terms that appear in the right hand side of (Dark Dimension and the Effective Field Theory limit), not only the first one. However, since this has no impact on the present discussion, we can equivalently refer to (10)1 or more completely to (Dark Dimension and the Effective Field Theory limit).
@@
Let us determine now the value of $\Lambda$ imposed by (11). First of all, we recall that from (4) we have $m_{{}_{\rm KK}}\sim\Lambda_{cc}^{1/4}$. A simple inspection of  (10)1 (or (Dark Dimension and the Effective Field Theory limit) with the replacement (11)) shows that this implies [^4] The same is true for (10)2.
@@
Eq. (20) definitely shows that what the authors of [3] do when they present our Eqs. (9) as Eqs. (10) (that are nothing but (9) with the replacement $\rho^{1l}\to\Lambda_{cc}$), is physically not allowed. They operate an inconsistent replacement, and Eqs. (10) are deprived of any physical meaning. The point is that the arguments they use in their attempt to confute our work completely depend on this replacement, actually cannot even be formulated without it. Their arguments are thus unfounded and inconsistent.
@@
The above analysis shows that Eqs. (10) are physically inconsistent, and obviously no use can be made of them. In particular, they cannot be used to confute our results. Unfortunately, this is precisely what the authors of [3] do, thus developing arguments and claims that are clearly deprived of any physical meaning.
The present work could end here. Nevertheless, in the hope that this can bring further clarity and help the reader gain more insight in the whole subject, we now provide a detailed analysis of their arguments, that (as already said) are based on the inconsistent Eqs. (10). To follow their presentation closely, below we consider both the SUSY and the non-SUSY cases, i.e. both their Eqs. (10)1 and (10)2.
@@
We finally comment on two additional arguments that the authors of [3] use to lend further support to their views. Making reference to the finiteness of string theory calculations, their goal is to give examples where finite results are obtained also in field theory. To this end, they consider (i) the calculation of the one-loop free energy $F^{1l}$ in finite temperature QFT, where the $T$-dependent contribution goes like $F_{T}^{1l}\sim T^{4}$ with no need for any “string regularization” (to use their words), and (ii) the calculation of the Casimir energy.
@@
For the case of KK theories ($\rho^{1l}$) the integer $n$ is related to the fifth component of the $(4+1)$D loop momentum $\hat{p}$, that is cut at $\Lambda$. Therefore $n$ and $p$ are intertwined between one another and are to be cut accordingly [1, 32]. This in particular implies an upper bound on the integer $n$. As a result, our Eqs. (25) and (26) of [1], here reported for the massless SUSY case in Eq. (Dark Dimension and the Effective Field Theory limit), are obtained. For the finite temperature field theory case ($F^{1l}$), on the contrary, the integer $n$ and the $3D$ momentum $p$ are not intertwined. The $d^{3}p$ integral provides the trace over quantum fluctuations (and is cut at a certain momentum cutoff), while the $\sum_{n}$ gives the statistical average (mixed states), and as such has to be performed all the way up to infinity to implement ergodicity. This infinite sum is at the origin of the finiteness [^9] It is worth to note that the full result for $F^{1l}$ is finite only if the theory is supersymmetric. of the T-dependent contribution $F_{T}^{1l}$ to $F^{1l}$. In fact,
@@
As mentioned above and discussed at length in [1] and [32], performing in the KK case too the sum over $n$ up to infinity (while keeping $p$ finite) mistreats the loop momentum. This is what causes the (artificial) disappearance of the $q$-dependent UV-sensitive terms in (Dark Dimension and the Effective Field Theory limit), giving the impression that there is a cancellation similar to what happens in the finite temperature case.
Moving now to the example of the Casimir energy $\mathcal{E}_{C}$, we observe that the latter is nothing but the difference between the vacuum energy $\rho_{{}_{R}}$ calculated in the KK theory with one compact dimension (that for our illustrative purposes we took as a circle of radius $R$) and the corresponding $\rho_{{}_{\infty}}$ calculated in the decompactification limit (i.e.  in the theory with all non-compact dimensions): $\mathcal{E}_{C}=\rho_{{}_{R}}-\rho_{{}_{\infty}}$. Now, if we perform the one-loop calculation of $\rho^{1l}$ according to previous literature, meaning that in (33) we perform the sum over $n$ up to infinity, $\mathcal{E}_{C}^{1l}$ takes a form in all similar to (Dark Dimension and the Effective Field Theory limit), with $T^{4}$ replaced by $m_{{}_{\rm KK}}^{4}$. However, we have shown in [1] that (contrary to what is found in previous literature) the subtraction $\rho_{{}_{R}}-\rho_{{}_{\infty}}$ is not in general sufficient to ensure the disappearance of all the UV-sensitive terms. This is for instance the case of KK theories with Scherk-Schwarz SUSY breaking considered in the present work.
@@
Before ending this Letter, we would like to make a couple of additional comments on some open questions connected with the subject discussed in the present work.
@@ Acknowledgments
I Appendix A - Dark Dimension scenario
@@
II Appendix B - Standard Model cutoff
@@
$\mathcal{S}^{(4)}_{\rm grav}=\frac{1}{2\kappa^{2}}\int\mathrm{d}^{4}x\,\sqrt{-g}$
×$\times\left[\mathcal{R}-2e^{2\alpha\phi}\hat{\Lambda}_{cc}+{2}\alpha\Box\phi+\frac{(\partial\phi)^{2}}{2}-\frac{e^{-6\alpha\phi}}{4}F^{2}\right],$ (43)
where the $4$D constant $\kappa=M_{P}^{2}$ is related to the $(4+1)$D $\hat{\kappa}=\hat{M}_{P}^{3}$ through the relation $\kappa^{2}=\hat{\kappa}^{2}/(2\pi R)$. The fields $\phi$ and $A_{\mu}$ in the above equation are dimensionless (dimensionful fields are obtained through the redefinition $\phi\to\phi/(\sqrt{2}\kappa),\,A_{\mu}\to A_{\mu}/(\sqrt{2}\kappa)$), and we used $02\alpha+\beta=$ 0. The canonical kinetic term in (II) for the radion field is obtained taking $\alpha={1}/{\sqrt{12}}$.
@@
$\mathcal{S}^{(4)}_{\rm mat}=\int d^{4}x\sqrt{-g}\,\,$ (44)
×$\times\sum_{n}\left[\left|D\varphi_{n}\right|^{2}-\left(e^{\sqrt{\frac{2}{3}}\frac{\phi}{M_{P}}}m^{2}+e^{\sqrt{6}\frac{\phi}{M_{P}}}\frac{n^{2}}{R^{2}}\right)\left|\varphi{{}_{n}}\right|^{2}\right],$
@@
where we defined $\Lambda_{\phi}\equiv\Lambda\,e^{\frac{1}{\sqrt{6}}\frac{\phi}{M_{P}}}$. In terms of the dimensionless $\phi$ of (42) and (II), and before using $\alpha=1/\sqrt{12}$, it is $\Lambda_{\phi}=e^{\alpha\phi}\Lambda=m_{{}_{\rm KK}}^{1/3}R^{1/3}\Lambda$.

### v4

@@ Abstract
In [1] we pointed out that in the Dark Dimension scenario [2] theoretical issues arise when the prediction for the vacuum energy $\rho$, that is obtained from swampland conjectures in string theory, is confronted with the corresponding result for $\rho$ in the effective field theory (EFT) limit. One of the problems concerns the widely spread belief that in higher dimensional EFTs with compact dimensions the vacuum energy is automatically finite. On the contrary, our analysis shows that $\rho$ contains (previously missed) UV-sensitive terms. Our work was challenged in [3]. Here we show why in our opinion the claims in [3] are flawed, and provide further support to our findings. We conclude presenting ideas on the physical mechanism that should dispose of the large UV contributions to $\rho$.
keywords:
dark energy; Kaluza-Klein theories; dark dimension.
{history}
1 Introduction
In search for a solution to the naturalness/hierarchy problem, in the nineties the possibility of having large extra dimensions (millimetre size) was considered [4, 5]. Recently, these ideas were applied to the so-called dark dimension (DD) scenario [2], according to which we might well live in a $(4+1)$D universe with one compact dimension of micrometre size. In a nutshell [^2] For convenience of the reader less familiar with the DD proposal, in Appendix A we provide a more detailed summary of the different aspects and steps that lead to this scenario. , the DD proposal [2] combines swampland conjectures with observational data and phenomenological bounds, and is based on the assumption that our universe lies in a unique (asymptotic) corner of the quantum gravity landscape, where only two kinds of light tower states are expected: towers of string excitations or towers of KK states. Indicating with $\mu_{tow}$ the tower scale, and with $\Lambda_{cc}$ the measured vacuum energy (cosmological constant times $M_{P}^{2}$, with $M_{P}$ the Planck mass), the application of the distance conjecture to (A)dS vacua relates the two as
@@ 1 Introduction
In a recent paper [1], we analysed aspects of this scenario, and pointed out possible theoretical issues. Below we briefly recall the scope and the main results of [1]. As already said, according to the proponents of the DD scenario, above the $m_{{}_{\rm KK}}$ scale our universe is described by a $(4+1)$D EFT with one compact dimension of micrometer size. This proposal has triggered a large amount of work, where $(4+1)$D EFT models that could realize the DD scenario are proposed, and their phenomenological consequences studied (see for instance [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]). In [1], our goal was to investigate on the matching between the string theory framework in which the basic relation (4) of the DD proposal is obtained and its low energy EFT limit.
@@
with no need for fine-tuning. In [1] we showed that the right hand side of (5) also contains UV-sensitive terms (see the first line in the right hand side of Eq. (2) below), that were missed in previous literature [^4] The presence of similar UV-sensitive terms in the calculation of the effective potential of a scalar theory was first hinted in [27]. Moreover, it was shown in [28] that the presence of Fayet-Iliopulos terms induces further divergences.. Therefore, the matching between the prediction (4) for the vacuum energy $\Lambda_{cc}$, that comes from swampland conjectures in string theory, and the corresponding result in the low energy field theory limit is not as straightforward as it was thought on the basis of the result (5). In this respect, it is worth to stress again (see the paragraph below Eq. (3)) that such a matching is crucial for the DD scenario.
In [1] we show why it is commonly thought that $\rho^{1l}$ is finite and proportional to $m_{{}_{\rm KK}}^{4}$: this result comes from a mistreatment of the loop momentum asymptotics. Performing the calculation properly, we found that (in addition to the finite term $\sim m^{4}_{{}_{\rm KK}}$ and to divergences that cancel out in SUSY theories) the one-loop contribution to the vacuum energy $\rho^{1l}$ contains “UV-sensitive” terms proportional to powers of the boundary charges. These divergences do not disappear, not even in a SUSY theory. This is due to the fact that, to trigger the Scherk-Schwarz mechanism, the boundary charges $q_{b}$ and $q_{f}$ of the boson and fermion superpartners need to be different.
Our work generated a debate that finally resulted in [3], where our results were challenged. Below we present a deep analysis of the arguments given in [3], and show why in our opinion these authors (in their attempt to confute our results) bring confusion on the whole subject using flawed arguments that lead them to incorrect conclusions. In view of the great wealth of theoretical and phenomenological work triggered by the DD scenario, we believe it worth trying to bring clarity on the subject and continuing its investigation. In the course of this analysis, we will bring new arguments that further support the conclusions of our previous work.
The rest of the paper is organised as follows. In section 2, we deeply investigate the arguments brought by the authors of [3] against our work, showing that they misuse our results and draw incorrect conclusions. Moreover, we present the global framework in which the DD scenario is framed, that allows to better understand the flaws contained in the claims made in [3]. In section 3, we show that arguments typically used in the literature (including [3]) to support the widely spread belief that in higher dimensional theories with compact dimensions the vacuum energy is automatically finite are misleading. In section 4, we present our conclusions, together with ideas on the physical mechanism that in our opinion should dispose of the large UV contributions to $\rho^{1l}$.
2 Global framework of the DD scenario and misuse of our results
As anticipated above, our results were challenged by the authors of [3]. Below we show that they misuse and distort our findings, and how this leads them to incorrect conclusions. Moreover, we present the global framework in which the DD scenario is framed (see Fig. 1), which further helps in bringing clarity on the whole subject. Before doing that, however, it is worth mentioning that [3] starts with the claim that we question the swampland relation (4), while what we did in [1] is totally different, and this was far from being our goal (see the Introduction above).
@@ 2 Global framework of the DD scenario and misuse of our results
As we will see, this result is important for our analysis. Let us also note that $m_{{}_{\rm KK}}$ is the cutoff of the “reduced 4D theory”, which besides brane fields contains the zero modes of bulk fields. Moreover, it is worth to observe that, although both the Standard Model and this “reduced theory” are $4$-dimensional EFTs, they are certainly not the same theory, and their cutoffs ($\Lambda_{\rm SM}$ and $m_{{}_{\rm KK}}$ respectively) should not be confused. In Fig. 1 we provide a comprehensive picture of the global framework in which the DD scenario is framed, described in detail in the caption below.
In Eq. (2), the divergences (i.e. UV-sensitive terms) are in the first line, the finite term in the second one. Limiting ourselves to consider only the most divergent term of $\rho^{1l}$ (and doing the same for the non-SUSY theory, whose result for $\rho^{1l}$ is not reported above but can be found in our work [1]), we have
@@
Although this is opposite to what we did (we repeat ourselves, in [1] our goal was to point out the presence of unexpected UV-sensitive terms), the replacement (11) is not in itself a problem. It is a theoretically legitimate possibility that the cutoff $\Lambda$ is a physical scale that appears in physical quantities, with no need for fine-tuning. However, the consequences of such a choice have to be fully taken into account. Anticipating on the results, we will show that the replacement $\rho^{1l}\to\Lambda_{cc}$ has crucial physical consequences that the authors of [3] failed to recognize. It is precisely this replacement that leads them to incorrect conclusions. Let us note that, for reasons that will be soon apparent, the right hand side of (10)1 should contain all the terms that appear in the right hand side of (2), not only the first one. However, since this has no impact on the present discussion, we can equivalently refer to (10)1 or more completely to (2).
@@
Let us determine now the value of $\Lambda$ imposed by (11). First of all, we recall that from (4) we have $m_{{}_{\rm KK}}\sim\Lambda_{cc}^{1/4}$. A simple inspection of  (10)1 (or (2) with the replacement (11)) shows that this implies [^5] The same is true for (10)2.
@@
Eq. (20) definitely shows that what the authors of [3] do when they present our Eqs. (9) as Eqs. (10) (that are nothing but (9) with the replacement $\rho^{1l}\to\Lambda_{cc}$), is physically not allowed. They operate an inconsistent replacement, and Eqs. (10) are deprived of physical meaning. The point is that the arguments they use in their attempt to confute our work completely depend on this replacement, actually cannot even be formulated without it. Their arguments are thus unfounded and inconsistent.
@@
The above analysis shows that Eqs. (10) are physically inconsistent, and obviously no use can be made of them. In particular, they cannot be used to confute our results. Unfortunately, this is precisely what the authors of [3] do, thus developing untenable arguments and claims.
The present section could end here. Nevertheless, in the hope that this can bring further clarity and help the reader gain more insight in the whole subject, we now provide a detailed analysis of their arguments, that (as already said) are based on the inconsistent Eqs. (10). To follow their presentation closely, below we consider both the SUSY and the non-SUSY cases, i.e. both their Eqs. (10)1 and (10)2.
@@
3 Casimir energy and finite temperature field theory
We finally comment on two additional arguments that the authors of [3] use to lend further support to their views101010 These arguments are not new, and have been repeatedly used over the years as they seem to substantiate the widely spread belief that for field theories with compact extra dimensions the vacuum energy turns out to be automatically finite.. Referring to the finiteness of string theory calculations, their goal is to give examples where finite results are obtained also in field theory. To this end, they consider (i) the calculation of the one-loop free energy $F^{1l}$ in finite temperature QFT, where the $T$-dependent contribution goes like $F_{T}^{1l}\sim T^{4}$ with no need for any “string regularization” (to use their words), and (ii) the calculation of the Casimir energy.
@@ 3 Casimir energy and finite temperature field theory
For the case of KK theories ($\rho^{1l}$) the integer $n$ is related to the fifth component of the $(4+1)$D loop momentum $\hat{p}$, that is cut at $\Lambda$. Therefore $n$ and $p$ are intertwined between one another and are to be cut accordingly [1, 32]. This in particular implies an upper bound on the integer $n$. As a result, our Eqs. (25) and (26) of [1], here reported for the massless SUSY case in Eq. (2), are obtained. For the finite temperature field theory case ($F^{1l}$), on the contrary, the integer $n$ and the $3D$ momentum $p$ are not intertwined. The $d^{3}p$ integral provides the trace over quantum fluctuations (and is cut at a certain momentum cutoff), while the $\sum_{n}$ gives the statistical average (mixed states), and as such has to be performed all the way up to infinity to implement ergodicity. This infinite sum is at the origin of the finiteness [^11] It is worth to note that the full result for $F^{1l}$ is finite only if the theory is supersymmetric. of the T-dependent contribution $F_{T}^{1l}$ to $F^{1l}$. In fact,
@@
As mentioned above and discussed at length in [1] and [32], performing in the KK case too the sum over $n$ up to infinity (while keeping $p$ finite) mistreats the loop momentum. This is what causes the (artificial) disappearance of the $q$-dependent UV-sensitive terms in (2), giving the impression that there is a cancellation similar to what happens in the finite temperature case.
Moving now to the example of the Casimir energy $\mathcal{E}_{C}$, we observe that the latter is nothing but the difference between the vacuum energy $\rho_{{}_{R}}$ calculated in the KK theory with one compact dimension (that for our illustrative purposes we took as a circle of radius $R$) and the corresponding $\rho_{{}_{\infty}}$ calculated in the decompactification limit (i.e.  in the theory with all non-compact dimensions): $\mathcal{E}_{C}=\rho_{{}_{R}}-\rho_{{}_{\infty}}$. Now, if we perform the one-loop calculation of $\rho^{1l}$ according to previous literature, meaning that in (33) we perform the sum over $n$ up to infinity, $\mathcal{E}_{C}^{1l}$ takes a form in all similar to (3), with $T^{4}$ replaced by $m_{{}_{\rm KK}}^{4}$. However, we have shown in [1] that (contrary to what is found in previous literature) the subtraction $\rho_{{}_{R}}-\rho_{{}_{\infty}}$ is not in general sufficient to ensure the disappearance of all the UV-sensitive terms. This is for instance the case of KK theories with Scherk-Schwarz SUSY breaking considered in the present work.
4 Summary, conclusions and open questions
@@
Before ending this work, we would like to make a couple of additional comments on some open questions connected with the subject discussed in the present work.
@@ Acknowledgments
Appendix A Dark Dimension scenario
@@
Appendix B Standard Model cutoff
@@
$\mathcal{S}^{(4)}_{\rm grav}=\frac{1}{2\kappa^{2}}\int\mathrm{d}^{4}x\,\sqrt{-g}\,\,\left[\mathcal{R}-2e^{2\alpha\phi}\hat{\Lambda}_{cc}+{2}\alpha\Box\phi+\frac{(\partial\phi)^{2}}{2}-\frac{e^{-6\alpha\phi}}{4}F^{2}\right],$ (43)
where the $4$D constant $\kappa=M_{P}^{2}$ is related to the $(4+1)$D $\hat{\kappa}=\hat{M}_{P}^{3}$ through the relation $\kappa^{2}=\hat{\kappa}^{2}/(2\pi R)$. The fields $\phi$ and $A_{\mu}$ in the above equation are dimensionless (dimensionful fields are obtained through the redefinition $\phi\to\phi/(\sqrt{2}\kappa),\,A_{\mu}\to A_{\mu}/(\sqrt{2}\kappa)$), and we used $02\alpha+\beta=$ 0. The canonical kinetic term in (43) for the radion field is obtained taking $\alpha={1}/{\sqrt{12}}$.
@@
$\mathcal{S}^{(4)}_{\rm mat}=\int d^{4}x\sqrt{-g}\,\,\sum_{n}\left[\left|D\varphi_{n}\right|^{2}-\left(e^{\sqrt{\frac{2}{3}}\frac{\phi}{M_{P}}}m^{2}+e^{\sqrt{6}\frac{\phi}{M_{P}}}\frac{n^{2}}{R^{2}}\right)\left|\varphi{{}_{n}}\right|^{2}\right],$ (44)
@@
where we defined $\Lambda_{\phi}\equiv\Lambda\,e^{\frac{1}{\sqrt{6}}\frac{\phi}{M_{P}}}$. In terms of the dimensionless $\phi$ of (42) and (43), and before using $\alpha=1/\sqrt{12}$, it is $\Lambda_{\phi}=e^{\alpha\phi}\Lambda=m_{{}_{\rm KK}}^{1/3}R^{1/3}\Lambda$.